# Negative Sampling in Recommendation: A Survey and Future Directions

### v1

@@ 2.2.2. Trade-off among Accuracy, Efficiency and Stability
The ”information overload” phenomenon has compelled recommenders to grapple with the issue of striking an appropriate trade-off among accuracy, efficiency, and stability (Chen et al., 2023a; Gao et al., 2022b). Negative sampling is no exception to this challenge (Ding et al., 2020; Li et al., 2021; Wang et al., 2020). Accuracy pertains to the precision of model’s predictions regarding user preferences, which is critical in evaluating the performance of recommenders. Efficiency focuses on how many queries the recommender can process within a unit of time, directly influencing its scalability and user experience. And stability concerns the consistent performance of the recommender across diverse datasets.
Nonetheless, these three aspects are generally interrelated in RS, making it challenging to balance them simultaneously. We furnish the sampling criteria, as well as the time complexities of candidate generation and negative selection for seven classic negative sampling algorithms, arranged in chronological order in Table 1. This illustrates that with the advancement in the accuracy of negative sampling techniques, there is a synchronous increase in their time complexity, underscoring the delicate trade-off between accuracy, efficiency and stability (Rendle et al., 2012; Rendle and Freudenthaler, 2014; Zhang et al., 2013; Ding et al., 2020; Shi et al., 2023; Huang et al., 2021; Ma et al., 2023). This is due to the fact that enhancing accuracy may lead to increased computational complexity, resulting in decreased efficiency. Some efficiency-improving optimization techniques may exhibit greater sensitivity to data variations and neglect specific details, potentially impacting the model’s accuracy and stability. Simultaneously, the pursuit of stability may increase time and space complexity, conflicting with the balance between accuracy and efficiency. Therefore, achieving the optimal balance among these three factors within negative sampling is a complex and challenging task.
@@ 2.2.3. Universality with Different Tasks, Objectives, and Datasets
In essence, the prevailing recommendation algorithms encompass different datasets with diverse distribution, distinct training objectives (e.g., matching and ranking), various recommendation scenarios (e.g., collaborative filtering, multi-modal recommendation, multi-behavior recommendation, sequence recommendation, and cross-domain recommendation), and diverse evaluation methods (random sampled testing, popularity-based sampled testing, and full testing). Although general negative sampling strategies (RNS (Rendle et al., 2012), PNS (Rendle and Freudenthaler, 2014), DNS (Zhang et al., 2013)) are available, some tailored negative sampling techniques designed for specific recommendation tasks can outperform these general methods (Cheng et al., 2021; Li et al., 2021; Liu et al., 2021b; Zhao et al., 2019; Shi et al., 2023; Gong et al., 2022; Zhang et al., 2013; Kim et al., 2021a; Yang et al., 2022a). Upon a comprehensive examination of the current negative sampling algorithms, it is evident that none of these approaches can perfectly match all recommendation tasks. In practical implementation, employing different negative sampling strategies for various tasks presents a challenge to the efficiency and stability of the recommender. This challenge underscores the obstacles that large-scale recommendation systems must overcome when transitioning to the real-world online deployment.
@@ 3. Literature review of Negative Sampling in Recommendation
Let and be the number of users and items respectively, we can define a user-item interaction matrix , where denotes the user has interacted with the item , and denotes has never interacted with . Negative sampling aims to select the appropriate instances from the unobserved item candidates as the negative samples to support the recommender optimization. We conduct a comprehensive literature review of existing negative sampling strategies in recommendation and classify them into five categories according to their technical trajectories: 1) Static negative sampling typically samples negative instances with the static probability; 2) Dynamic negative sampling attempts to dynamically select negative instances with the pre-established sampling criterion; 3) Adversarial negative generation leverages the adversarial learning paradigm to sample or generate plausible items as the negative samples; 4) Importance re-weighting aims to identify the importance of each sample and assign diverse weights for them in a data-driven manner; and 5) Knowledge-enhanced negative sampling focuses on sampling negative instances with the mining of the implicit association from the auxiliary knowledge. In the following, we detail the effectiveness mechanisms and meaningful insights of these negative sampling strategies.
@@ 3.1. Static Negative Sampling
Traditional approaches in the initial stages of deep learning for recommendation relied on the static negative sampling (SNS) method whereby several items are selected from users’ unobserved items. The primary objective of SNS is to offer a diverse range of informative negative samples and facilitate the acquisition of more comprehensive user preference patterns by the recommender. We delineate existing SNS studies into four categories, Uniform SNS, Predefined SNS, Popularity-based SNS, and Non-sampling SNS. The in-depth investigation of behavior dependencies, advantages, and challenges of them are documented in Table 2. Among them, Uniform SNS emerges as the most widely employed SNS method with its easy-to-deploy and time-efficient character. However, the randomness inherent in sampling introduces variability to its recommendation performance. Predefined SNS leverages authentic user-centric ratings for negative sample selection. Despite its straightforward implementation, it is critically reliant on the accessibility of user behaviors. In the incorporation of the mainstream preferences of the majority of users into the training process, Popularity-based SNS also introduces the inherent popularity bias and conformity bias. Non-sampling SNS incorporates the unobserved items of each user from the whole training data. Maintaining visibility over the entire corpus of the dataset enhances the recommender’s performance, yet encountering challenges in computational efficiency.
@@ 3.1.1. Uniform SNS
Uniform SNS randomly selects items from the unobserved interactions of the user as the negative sample (Zhou et al., 2018). This strategy is extensively employed across diverse recommendation scenarios (e.g., Collaborative Filtering (Rendle et al., 2012; He et al., 2017), Graph-based Recommendaiton (Wang et al., 2019a; He et al., 2023a), Sequential Recommendaiton (Yang et al., 2022b; Zou et al., 2024), Multi-modal Recommendation (Liu et al., 2019; Wang et al., 2023a, 2024), and Cross-domain Recommendation (Gao et al., 2021a; Cao et al., 2022)) and distinct recommendation applications (e.g., Point-of-Interest Recommendation (Yang et al., 2017; Lian et al., 2014), News Recommendation (Wu et al., 2019d, c), Song Recommendation (Tran et al., 2019) and API Recommendation (Wei et al., 2022)). Primarily, its ease-to-deploy mechanism makes it well-suited for large-scale RS, as it obviates the necessity for additional computations. Additionally, it introduces a degree of diversity, mitigating the risk of over-fitting and consequently enhancing recommender’s generalization and robustness. Moreover, this strategy facilitates the exploration of items that users have not interacted with, thereby aiding the recommender in uncovering novel recommendation opportunities and unexplored areas of user preference.
@@
| Uniform SNS (Rendle et al., 2012; Petrov and Macdonald, 2022; Wang et al., 2023a, 2024; Zou et al., 2024) |
@@
| Predefined SNS (Song et al., 2015; Yu et al., 2018; Zheng et al., 2018; Lin et al., 2024; Xi et al., 2023; Dai et al., 2023; Li et al., 2023) | User-centric rating |
@@
| Popularity-based SNS (Gantner et al., 2012; Cheng et al., 2021; Togashi et al., 2021; Wang et al., 2019b; Li et al., 2018; He et al., 2016) |
@@
It is noteworthy that TDM (Zhu et al., 2018) has made advancements in enhancing the uniform SNS method within the hierarchical tree structure. It denotes the positive sample set as the leaf node that has interacted with the user and their ancestor nodes, while randomly selecting nodes except positive ones in each level to constitute the set of negative samples. RSS (Petrov and Macdonald, 2022) also tries to propose a recency-based positive sampling strategy to generate multiple positive training samples out of a single user sequence simultaneously, which is orthogonal to negative sampling and can be applied independently. However, given that the primary focus of this subsection lies in the examination of the negative sampling strategies in recommendation, there has been limited elaboration on the positive-augmented sampling methods.
3.1.2. Predefined SNS
Predefined SNS incorporates the pre-defined negative samples from the dataset into the recommenders’s training process (Song et al., 2015; Yu et al., 2018; Sun et al., 2021; Zhang et al., 2024). For example, DRN (Zheng et al., 2018) employs real user behaviors (e.g., skipped, clicked, and ordered actions) within the dataset to delineate positive and negative samples. Additionally, a series of studies designate the ratings of in the explicit user interactions as positive samples. Within this context, ELLR (Song et al., 2015) and IF-BPR (Yu et al., 2018) segregate ratings of as negative samples, while SCoRe (Qin et al., 2020), ML-MGC (Sun et al., 2021) and many LLM-based recommenders (Lin et al., 2024; Xi et al., 2023; Dai et al., 2023; Li et al., 2023) extend this categorization to include ratings of as negative samples.
3.1.3. Popularity-based SNS
Popularity-based SNS selects negative samples based on the popularity of items, that is, the more popular the item is, the more likely it is to be selected as the negative sample (Gantner et al., 2012; Quadrana et al., 2017; Cheng et al., 2021; Ma et al., 2018; Rendle and Freudenthaler, 2014; Togashi et al., 2021; Wang et al., 2019b; Li et al., 2018; He et al., 2016). Relying on the assumption that the popularity of items may demonstrate users’ global preferences, a series of popularity-based SNS methods (e.g., WBPR (Gantner et al., 2012), HRNN (Quadrana et al., 2017), NNCF (Rendle and Freudenthaler, 2014), BINN (Li et al., 2018)) typically assign sampling weights to items based on their frequency. This encourages the recommender to incorporate more popular items as negative samples into the training process. For quick updating the recommender parameters with the new incoming data, eALS (He et al., 2016) assigns the weight of missing data based on the popularity of items by an incremental update strategy. SoftRec (Cheng et al., 2021) allows some unobserved items to have non-zero supervisory signals by generating soft targets with popularity distribution, thereby incorporating more unobserved knowledge into the training process. Similar to RSS (Petrov and Macdonald, 2022), SAE-NAD (Ma et al., 2018) adopts the general weighting scheme to assign higher confidence to each positive sample based on its frequency. In contrast, it assigns the weights of all negative examples to the same value (e.g., 1.0) for the Point-of-Interest recommendation task.
Notably, KGPL (Togashi et al., 2021) and MKR (Wang et al., 2019b) employ the Popularity-based SNS for the graph structure. In this context, KGPL (Togashi et al., 2021) directly formulates the sampling confidence of negative samples based on the interaction frequencies of each item across the entire dataset. Furthermore, it assigns reliable pseudo-labels to the interacted items of cold-start users by considering items that can be reached through meta-paths (i.e., the paths rooted at a user’s interacted items in KG). Different from the above studies, MKR (Wang et al., 2019b) follows the negative sampling strategy of the classical work (Mikolov et al., 2013) in NLP, which reduces the sampling probability for the high-frequency items and augments it for the low-frequency items. This is motivated by the observation that less probable words tend to be more pivotal, while higher probability words such as ‘the’, ‘a’, and ‘to’ bear less weight in NLP tasks. It actually deviates from the core assumptions in recommendation, thus MKR adopts such a sampling strategy primarily to enhance the computational efficiency.
3.1.4. Non-sampling SNS
Non-sampling SNS considers the unobserved instances from the whole training data for recommender learning, thus avoiding negative sampling (Chen et al., 2020d, c, 2019d, b; Li et al., 2021). These related works argue that negative sampling strategies are highly sensitive to the data distribution and the number of negative samples, making them difficult to achieve the optimal performance in large-scale RS. To converge to a better optimum, EATNN (Chen et al., 2019d) first devises an efficient optimization method to learn from the whole training set without negative sampling, and reformulates a square loss function with rigorous mathematical reasoning for accelerating the training process in social-aware recommendation. ENMF (Chen et al., 2020c) stands as an extended edition of EATNN (Chen et al., 2019d) that extends its efficient learning method with the item-based and alternating-based forms, enabling its application for implicit feedback. Due to the behavior characteristic in multi-bahevior recommendation, EHCF (Chen et al., 2020d) links the prediction of each type of behavior in a transfer manner and applies non-sampling optimization for the recommender by considering all samples in each parameter update. Inspired by the memorization strategies, ENSFM (Chen et al., 2020b) employs the non-sampling technique to maintain the convergence into a better optimum in a stable way. JNSKR (Li et al., 2021) refine the existing non-sampling strategy into the modeling of the knowledge graph embeddings, which consists of entity-relation-entity triplets. Furthermore, it devises an efficient optimization approach to learn more precise entity embeddings and item representations within the knowledge graph, resulting in a discernible and stable enhancement of recommendation performance.

### v2

@@ 2.4.2. Trade-off among Accuracy, Efficiency and Stability
The ”information overload” phenomenon has compelled recommenders to grapple with the issue of striking an appropriate trade-off among accuracy, efficiency, and stability (Chen et al., 2023a; Gao et al., 2024). Negative sampling is no exception to this challenge (Ding et al., 2020; Li et al., 2021; Wang et al., 2020). Accuracy pertains to the precision of model’s predictions regarding user preferences, which is critical in evaluating the performance of recommenders. Efficiency focuses on how many queries the recommender can process within a unit of time, directly influencing its scalability and user experience. And stability concerns the consistent performance of the recommender across diverse datasets.
Nonetheless, these three aspects are generally interrelated in RS, making it challenging to balance them simultaneously. We furnish the sampling criteria, as well as the time complexities of candidate generation and negative selection for seven classic negative sampling algorithms, arranged in chronological order in Table 1. This illustrates that with the advancement in the accuracy of negative sampling techniques, there is a synchronous increase in their time complexity, underscoring the delicate trade-off between accuracy, efficiency and stability (Rendle et al., 2009; Rendle and Freudenthaler, 2014; Zhang et al., 2013; Ding et al., 2020; Shi et al., 2023; Huang et al., 2021; Ma et al., 2023). This is due to the fact that enhancing accuracy may lead to increased computational complexity, resulting in decreased efficiency. Some efficiency-improving optimization techniques may exhibit greater sensitivity to data variations and neglect specific details, potentially impacting the model’s accuracy and stability. Simultaneously, the pursuit of stability may increase time and space complexity, conflicting with the balance between accuracy and efficiency. Therefore, achieving the optimal balance among these three factors within negative sampling is a complex and challenging task.
@@ 2.4.3. Universality with Different Tasks, Objectives, and Datasets
In essence, the prevailing recommendation algorithms encompass different datasets with diverse distribution, distinct training objectives (e.g., matching and ranking), various recommendation scenarios (e.g., collaborative filtering, multi-modal recommendation, multi-behavior recommendation, sequence recommendation, and cross-domain recommendation), and diverse evaluation methods (random sampled testing, popularity-based sampled testing, and full testing). Although general negative sampling strategies (RNS (Rendle et al., 2009), PNS (Rendle and Freudenthaler, 2014), DNS (Zhang et al., 2013)) are available, some tailored negative sampling techniques designed for specific recommendation tasks can outperform these general methods (Cheng et al., 2021; Li et al., 2021; Liu et al., 2021b; Zhao et al., 2019; Shi et al., 2023; Gong et al., 2022; Zhang et al., 2013; Kim et al., 2021a; Yang et al., 2022a). Upon a comprehensive examination of the current negative sampling algorithms, it is evident that none of these approaches can perfectly match all recommendation tasks. In practical implementation, employing different negative sampling strategies for various tasks presents a challenge to the efficiency and stability of the recommender. This challenge underscores the obstacles that large-scale recommender systems must overcome when transitioning to the real-world online deployment.
@@ 3. Literature review of Negative Sampling in Recommendation
Let and be the number of users and items , we can define a user-item interaction matrix , where denotes the user has interacted with the item , and denotes has never interacted with . Negative sampling aims to select the appropriate instances from the unobserved item candidates as the negative samples to support the recommender optimization. We conduct a comprehensive literature review of existing negative sampling strategies in recommendation, classify them into five categories according to their technical trajectories, provide the overall ontology of our taxonomy in Fig. 4 and additionally offer five tables (from Table 2 to Table 6) to illustrate their core contributions, strengths, and drawbacks: 1) Static negative sampling typically samples negative instances with the static probability; 2) Dynamic negative sampling attempts to dynamically select negative instances with the pre-established sampling criterion; 3) Adversarial negative generation leverages the adversarial learning paradigm to sample or generate plausible items as the negative samples; 4) Importance re-weighting aims to identify the importance of each sample and assign diverse weights for them in a data-driven manner; and 5) Knowledge-enhanced negative sampling focuses on sampling negative instances with the mining of the implicit association from the auxiliary knowledge. In the following, we detail the effectiveness mechanisms and meaningful insights of these negative sampling strategies.
@@ 3.1. Static Negative Sampling
Traditional approaches in the initial stages of deep learning for recommendation relied on the static negative sampling (SNS) method whereby several items are selected from users’ unobserved items. The primary objective of SNS is to offer a diverse range of informative negative samples and facilitate the acquisition of more comprehensive user preference patterns by the recommender. We will introduce several types of SNS strategies (Uniform SNS, Predefined SNS, Popularity-based SNS, and Non-sampling SNS) in the following subsections.
@@ 3.1.1. Uniform SNS
Uniform SNS randomly selects items from the unobserved interactions of the user as the negative sample (Zhou et al., 2018; Dai et al., 2024a). This strategy is extensively employed across diverse recommendation scenarios (e.g., Collaborative-guided Recommendation (Rendle et al., 2009; He et al., 2017, 2023a; Wu et al., 2023b), Sequential Recommendaiton (Yang et al., 2022b; Zou et al., 2024), Social Recommendation (Yang et al., 2023c, 2024; Wu et al., 2023a, c), Multi-modal Recommendation (Wang et al., 2023a, 2024a), and Cross-domain Recommendation (Cao et al., 2022; Wang et al., 2024b)) and distinct recommendation applications (e.g., Point-of-Interest Recommendation (Lian et al., 2014), News Recommendation (Wu et al., 2019c), Song Recommendation (Tran et al., 2019) and API Recommendation (Wei et al., 2022)). Primarily, its ease-to-deploy mechanism makes it well-suited for large-scale RS, as it obviates the necessity for additional computations. It also introduces a degree of diversity, mitigating the risk of over-fitting and consequently enhancing recommender’s generalization and robustness. Moreover, this strategy facilitates the exploration of items that users have not interacted with, thereby aiding the recommender in uncovering novel recommendation opportunities and unexplored areas of user preference.
Notably, TDM (Zhu et al., 2018) denotes the positive sample as the leaf nodes that have interacted with the user and their ancestor nodes within the hierarchical tree structure, while randomly selecting nodes except positive ones in each level to constitute the set of negative samples. RSS (Petrov and Macdonald, 2022) also tries to propose a recency-based strategy to generate multiple positive samples out of a single user sequence simultaneously, which is orthogonal to negative sampling and can be applied independently. However, given that the primary focus of this survey lies in the examination of negative sampling in recommendation, there has been limited elaboration on the positive-augmented sampling methods.
3.1.2. Predefined SNS
Predefined SNS incorporates the pre-defined negative samples from the dataset into the recommenders’s training process (Song et al., 2015; Yu et al., 2018; Sun et al., 2021; Zhang et al., 2024b). For example, DRN (Zheng et al., 2018) employs real user behaviors (e.g., skipped, clicked, and ordered actions) within the dataset to delineate positive and negative samples. Additionally, a series of studies designate the ratings of in the explicit user interactions as positive samples. Within this context, ELLR (Song et al., 2015) and IF-BPR (Yu et al., 2018) segregate ratings of as negative samples, while SCoRe (Qin et al., 2020), ML-MGC (Sun et al., 2021) and many LLM-based recommenders (Lin et al., 2024; Xi et al., 2024; Dai et al., 2023; Li et al., 2023) extend this categorization to include ratings of as negative samples.
3.1.3. Popularity-based SNS
Relying on the assumption that the popularity of items may demonstrate users’ global preferences in the specific domain, popularity-based SNS typically selects negative samples based on the popularity of items, that is, the more popular the item is, the more likely it is to be selected as the negative sample (Gantner et al., 2012; Cheng et al., 2021; Ma et al., 2018; Rendle and Freudenthaler, 2014; Togashi et al., 2021; Wang et al., 2019; Li et al., 2018; He et al., 2016; Yang et al., 2023d). This encourages the recommender to incorporate more popular items as negative samples into the training process. For quick updating the recommender parameters with the new incoming data, eALS (He et al., 2016) assigns the weight of missing data based on the popularity of items by an incremental update strategy. SoftRec (Cheng et al., 2021) allows some unobserved items to have non-zero supervisory signals by generating soft targets with popularity distribution, thereby incorporating more unobserved knowledge into the training process. Similar to RSS (Petrov and Macdonald, 2022), SAE-NAD (Ma et al., 2018) adopts the general weighting scheme to assign higher confidence to each positive sample based on its frequency. In contrast, it assigns the weights of all negative examples to the same value (e.g., 1.0) for the Point-of-Interest recommendation task.
Notably, KGPL (Togashi et al., 2021) and MKR (Wang et al., 2019) employ the Popularity-based SNS on the graph structure. For instance, KGPL (Togashi et al., 2021) directly formulates the sampling confidence of negative samples based on their interaction frequencies across the entire dataset. Furthermore, it assigns reliable pseudo-labels to the interacted items of cold-start users by considering the reachable items under given meta-paths (i.e., the paths rooted at a user’s interacted items in KG). Different from the above studies, MKR (Wang et al., 2019) follows the negative sampling strategy of the classical work (Mikolov et al., 2013) in NLP, which reduces the sampling probability for the high-frequency items and augments it for the low-frequency items. This is motivated by the observation that less probable words tend to be more pivotal, while higher probability words such as ‘the’, ‘a’, and ‘to’ bear less weight in NLP tasks. It actually deviates from the core assumptions in recommendation, thus MKR adopts such a sampling strategy primarily to enhance the computational efficiency.
3.1.4. Non-sampling SNS
Non-sampling SNS considers the unobserved instances from the whole training data for recommender learning, thus avoiding negative sampling (Chen et al., 2020d, c, 2019d, b; Li et al., 2021). This strategy argues that negative sampling are highly sensitive to the data distribution and the number of negative samples, making them difficult to achieve the optimal performance in large-scale RS. To converge to a better optimum, EATNN (Chen et al., 2019d) first devises an efficient optimization method to learn from the whole training set without negative sampling, and reformulates a square loss function with rigorous mathematical reasoning for accelerating the training process in social-aware recommendation. ENMF (Chen et al., 2020c) stands as an extended edition of EATNN (Chen et al., 2019d) that extends its efficient learning method with the item-based and alternating-based forms, enabling its application for implicit feedback. Due to the behavior characteristic in multi-bahevior recommendation, EHCF (Chen et al., 2020d) links the prediction of each type of behavior in a transfer manner and applies non-sampling optimization for the recommender by considering all samples in each parameter update. Inspired by the memorization strategies, ENSFM (Chen et al., 2020b) employs the non-sampling technique to maintain the convergence into a better optimum in a stable way. JNSKR (Li et al., 2021) refine the existing non-sampling strategy into the modeling of the knowledge graph embeddings, which consists of entity-relation-entity triplets. Furthermore, it devises an efficient optimization approach to learn more precise entity embeddings and item representations within the knowledge graph, resulting in a stable enhancement of recommendation performance.
@@ 3.1.4. Non-sampling SNS
| Uniform SNS (Rendle et al., 2009; Petrov and Macdonald, 2022; Wang et al., 2023a, 2024a; Zou et al., 2024) |
@@
| Predefined SNS (Song et al., 2015; Yu et al., 2018; Zheng et al., 2018; Lin et al., 2024; Xi et al., 2024; Dai et al., 2023; Li et al., 2023) | User-centric rating |
@@
| Popularity-based SNS (Gantner et al., 2012; Cheng et al., 2021; Togashi et al., 2021; Wang et al., 2019; Li et al., 2018; He et al., 2016) |
@@
Summary: The in-depth investigation of behavior dependencies, advantages, and challenges of these above four SNS categories are documented in Table 2. Among them, Uniform SNS emerges as the most widely employed SNS method with its easy-to-deploy and time-efficient character. However, the randomness inherent in sampling introduces variability to its recommendation performance. Predefined SNS leverages authentic user-centric ratings for negative sample selection. Despite its straightforward implementation, it is critically reliant on the accessibility of user behaviors. In the incorporation of the mainstream preferences of the majority of users into the training process, Popularity-based SNS also introduces the inherent popularity bias and conformity bias. Non-sampling SNS incorporates the unobserved items of each user from the whole training data. Maintaining visibility over the entire corpus of the dataset enhances the recommender’s performance, yet encountering challenges in computational efficiency.